# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

### v7

@@ 4 Homomorphic Training
Adagrad method is not a practical solution for homomorphic LR due to its frequent inversion operations. It seems plausible that the enhanced NAG is likely the reasonable choice for homomorphic LR training. We adopt the enhanced NAG method to implement secure logistic regression training based on HE. The difficulty in applying the quadratic gradient is to invert the diagonal matrix $\tilde{B}$ in order to obtain $\bar{B}$. We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud. Since data owner has to prepare the dataset and normalize it, it would also be practicable for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
@@
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. (2018a), utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
@@
Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al. (2018a) used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$. The weight vector $\beta^{(0)}$ consisting of zeros and the diagnoal elements of $\bar{B}$ are copied $n$ times to form two matrices. The data owner then encrypt the two matrices into two ciphertexts $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$, respectively. The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
@@
X $=\left[\begin{array}[]{cccc}1&x_{11}&\ldots&x_{1d}\\ 1&x_{21}&\ldots&x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n1}&\ldots&x_{nd}\\ \end{array}\right],\text{Y}=\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\\ \end{array}\right],$
$\text{ct}_{Z}$ $=Enc\left[\begin{array}[]{cccc}y_{1}&y_{1}x_{11}&\ldots&y_{1}x_{1d}\\ y_{2}&y_{2}x_{21}&\ldots&y_{2}x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ y_{n}&y_{n}x_{n1}&\ldots&y_{n}x_{nd}\\ \end{array}\right],$
@@
$\text{ct}_{\beta}^{(0)}$ $=Enc\left[\begin{array}[]{cccc}\beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \vdots&\vdots&\ddots&\vdots\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \end{array}\right],$
$\text{ct}_{\bar{B}}$ $=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
@@
Edinburgh 1253 9 Ours 5 4 0.04 0.6 89.52 0.943
(Kim et al., 2018a) 5 7 0.02 3.6 91.04 0.958
lbw 189 9 Ours 5 4 0.04 0.6 71.35 0.667
(Kim et al., 2018a) 5 7 0.02 3.3 69.19 0.689
nhanes3 15649 15 Ours 5 4 0.31 4.5 79.23 0.637
(Kim et al., 2018a) 5 7 0.16 7.3 79.22 0.717
pcs 379 9 Ours 5 4 0.04 0.6 63.20 0.733
(Kim et al., 2018a) 5 7 0.02 3.5 68.27 0.740
uis 575 8 Ours 5 4 0.04 0.6 74.43 0.597
(Kim et al., 2018a) 5 7 0.02 3.5 74.44 0.603
@@ 5 Experiments
For a fair comparison with (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets. Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators. Tables 1 and 2 display the results of the two experiments, respectively. The two tables also provide the average evaluation running time for each iteration and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method). We adopt the same packing method that Kim et al. (2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$. We chose $1+0.9^{t}$ as our learning rate configuration.
@@ 6 Conclusion
The quadratic gradient presented in this work can be constructed from the Hessian matrix directly, and thus somehow combines the first-order Newton’s method and the second-order gradient (descent) method together. There is a promising chance that quadratic gradient could accelerate other gradient methods, which is an open future work.
@@
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v8

@@ 4 Secure Training
We adopt the enhanced NAG method to implement secure logistic regression training based on HE. The difficulty in applying the quadratic gradient is to invert the diagonal matrix $\tilde{B}$ in order to obtain $\bar{B}$. We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud. Since data owner has to prepare the dataset and normalize it, it would also be practicable for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
@@
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. Kim et al. (2018a), utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
@@
Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al. Kim et al. (2018a) used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$. The weight vector $\beta^{(0)}$ consisting of zeros and the diagnoal elements of $\bar{B}$ are copied $n$ times to form two matrices. The data owner then encrypt the two matrices into two ciphertexts $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$, respectively. The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
@@
$\text{X}=\left[\begin{array}[]{cccc}1&x_{11}&\ldots&x_{1d}\\ 1&x_{21}&\ldots&x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n1}&\ldots&x_{nd}\\ \end{array}\right],\text{Y}=\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\\ \end{array}\right],$ $\text{ct}_{Z}=Enc\left[\begin{array}[]{cccc}y_{1}&y_{1}x_{11}&\ldots&y_{1}x_{1d}\\ y_{2}&y_{2}x_{21}&\ldots&y_{2}x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ y_{n}&y_{n}x_{n1}&\ldots&y_{n}x_{nd}\\ \end{array}\right],$
@@
$\text{ct}_{\beta}^{(0)}=Enc\left[\begin{array}[]{cccc}\beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \vdots&\vdots&\ddots&\vdots\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \end{array}\right],$ $\text{ct}_{\bar{B}}=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
@@
Edin 1253 8 Ours 4 0.04 0.6 89.52 0.943
Baseline 7 0.02 3.6 91.04 0.958
lbw 189 8 Ours 4 0.04 0.6 71.35 0.667
Baseline 7 0.02 3.3 69.19 0.689
nhanes3 15649 15 Ours 4 0.31 4.5 79.23 0.637
Baseline 7 0.16 7.3 79.22 0.717
pcs 379 9 Ours 4 0.04 0.6 63.20 0.733
Baseline 7 0.02 3.5 68.27 0.740
uis 575 8 Ours 4 0.04 0.6 74.43 0.597
Baseline 7 0.02 3.5 74.44 0.603
@@ 5 Experiments
For a fair comparison with the baseline (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets. Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators. Tables 1 and 2 display the results of the two experiments, respectively. The two tables also provide the average evaluation running time for each iteration and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method). We adopt the same packing method that Kim et al. Kim et al. (2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$. We chose $1+0.9^{t}$ as our learning rate configuration.
@@ 6 Conclusion
The quadratic gradient presented in this work can be constructed from the Hessian matrix directly, and thus somehow combines the first-order Newton’s method and the second-order gradient (descent) method together. There is a promising chance that quadratic gradient could accelerate other gradient methods.
@@
It is highly likely that the quadratic gradient approach could combine first-order gradient descent/ascent algorithms with the second-order Newtonian method, and that it could be applicable to a wide range of numerical optimization problems.